OpenStudy (anonymous):

Point X is located at (2, −6) and point Z is located at (0, 5). Find the y value for the point Y that is located one over five the distance from point X to point Z. −3.6 −3.8 −4 −4.2

OpenStudy (anonymous):

@navk

OpenStudy (anonymous):

sorry I had signed out. saw this now Again using section formula, y = \[y = \frac{1 \times 5 + 5 \times -6}{1 + 5}\] Solving it, it gives -25/6 = -4.2 (I answered the same one earlier yesterday for you)

OpenStudy (anonymous):

Is the answer not matching?